# Two-Timescale Gradient Descent Ascent Algorithms for Nonconvex Minimax Optimization

### v1

@@ Abstract
We provide a unified analysis of two-timescale gradient descent ascent (TTGDA) for solving structured nonconvex minimax optimization problems in the form of $\min_{\mathbf{x}}\max_{\mathbf{y}\in\mathcal{Y}}f(\mathbf{x},\mathbf{y})$, where the objective function $f(\mathbf{x},\mathbf{y})$ is nonconvex in $\mathbf{x}$ and concave in $\mathbf{y}$, and the constraint set $\mathcal{Y}\subseteq\mathbb{R}^{n}$ is convex and bounded. In the convex-concave setting, the single-timescale GDA achieves strong convergence guarantees and has been used for solving application problems arising from operations research and computer science. However, it can fail to converge in more general settings. Our contribution in this paper is to design the simple deterministic and stochastic TTGDA algorithms that efficiently find one stationary point of the function $\Phi(\cdot):=\max_{\mathbf{y}\in\mathcal{Y}}f(\cdot,\mathbf{y})$. Specifically, we prove the theoretical bounds on the complexity of solving both smooth and nonsmooth nonconvex-concave minimax optimization problems. To our knowledge, this is the first systematic analysis of TTGDA for nonconvex minimax optimization, shedding light on its superior performance in training generative adversarial networks (GANs) and in solving other real-world application problems.
@@ 1 Introduction
In the wake of von Neumann’s seminal work (Neumann, 1928), the problem of finding a saddle point solution to minimax optimization problems, under various assumptions on the objective functions, has been a major focus of research in many areas, including economics, control theory, operations research and computer science (Basar and Olsder, 1999; Nisan et al., 2007; Von Neumann and Morgenstern, 2007). Recently, the minimax optimization theory has begun to see applications in machine learning, with examples including generative adversarial networks (GANs) (Goodfellow et al., 2014), robust statistics (Xu et al., 2009; Abadeh et al., 2015), online adversarial learning (Cesa-Bianchi and Lugosi, 2006), robust training of deep neural networks (Sinha et al., 2018) and distributed computing over networks (Shamma, 2008; Mateos et al., 2010). Current trends in multi-agent learning with large-scale economic systems seem likely to further increase the need for efficient algorithms for computing minimax optima and related equilibrium solutions.
@@
Formally, we present the mathematical form of generic minimax optimization problem as follows,
@@
where $f:\mathbb{R}^{m}\times\mathbb{R}^{n}\mapsto\mathbb{R}$ is the objective function and $\mathcal{Y}$ is a constraint set. One of simplest methods for solving the minimax optimization problems is to generalize gradient descent (GD) so that each loop consists in both a descent step and an ascent step. The algorithm, referred to as gradient descent ascent (GDA), performs one-step gradient descent over the min variable $\mathbf{x}$ with a learning rate $\eta_{\mathbf{x}}>0$ and one-step gradient ascent over the max variable $\mathbf{y}$ with another learning rate $\eta_{\mathbf{y}}>0$.
On the positive side, when $f$ is convex in $\mathbf{x}$ and concave in $\mathbf{y}$, there is a vast amount of literature on asymptotic and nonasymptotic convergence for average iterates generated by classical single-timescale GDA (i.e., $\eta_{\mathbf{x}}=\eta_{\mathbf{y}}$(see, e.g., Korpelevich, 1976; Chen and Rockafellar, 1997; Nemirovski, 2004; Auslender and Teboulle, 2009; Nedić and Ozdaglar, 2009). In addition, the local linear convergence can be established under an additional assumption that $f$ is locally strongly convex in $\mathbf{x}$ and locally strongly concave in $\mathbf{y}$ (Cherukuri et al., 2017; Adolphs et al., 2019). In contrast, there has been no shortage of negative results that highlight the fact that in the general settings, the single-timescale GDA might converge to limit cycles or even diverge (Mertikopoulos et al., 2018).
@@
We provide detailed pseudocode for TTGDA and its stochastic variant (TTSGDA) in Algorithms 1 and 2. Note that Algorithm 2 repeatedly samples while remaining at the current iterate and the subroutine $(\hat{\bf g}_{\mathbf{x}}^{t},\hat{\bf g}_{\mathbf{y}}^{t})=\textsf{SG}(G,\{\xi_{i}^{t}\}_{i=1}^{M},\mathbf{x}_{t},\mathbf{y}_{t})$ can be implemented in general by the updates
$\hat{\bf g}_{\mathbf{x}}^{t}=\tfrac{1}{M}\left(\sum_{i=1}^{M}G_{\mathbf{x}}(\mathbf{x}_{t},\mathbf{y}_{t},\xi_{i}^{t})\right),\quad\hat{\bf g}_{\mathbf{y}}^{t}=\tfrac{1}{M}\left(\sum_{i=1}^{M}G_{\mathbf{y}}(\mathbf{x}_{t},\mathbf{y}_{t},\xi_{i}^{t})\right),$
@@
Our main results in this paper are presented in four theorems. In particular, the first and second theorems (Theorems 4.4 and 4.6) provide a complete characterization of TTGDA and TTSGDA when applied to solve smooth minimax optimization problems. In the nonconvex-strongly-concave setting, TTGDA and TTSGDA require $O(\kappa^{2}\epsilon^{-2})$ and $O(\kappa^{3}\epsilon^{-4})$ (stochastic) gradient evaluations to find an $\epsilon$-stationary point (cf. Definition 3.3) where $\kappa$ is a condition number. In the nonconvex-concave setting, TTGDA and TTSGDA require $O(\epsilon^{-6})$ and $O(\epsilon^{-8})$ (stochastic) gradient evaluations to find an $\epsilon$-stationary point (cf. Definition 3.8). These results have appeared in our preliminary work (Lin et al., 2020a) which was accepted by ICML. However, our analysis here is simpler and more elegant.
@@
Another two theorems (Theorems 5.4 and 5.6) are more substantive contributions of this paper. In particular, our theorems show that TTGDA and TTSGDA can efficiently find one $\epsilon$-stationary point (cf. Definition 3.8) even when applied to solve nonsmooth minimax optimization problems. The complexity bounds of $\tilde{O}(\epsilon^{-6})$ and $O(\epsilon^{-8})$ in terms of (stochastic) gradient evaluations can be established for the nonconvex-strongly-concave and nonconvex-concave settings. These results are strong since the aforementioned machine learning applications often have the nonsmooth objective functions (e.g., the use of ReLU neural network in GANs). Our theorems demonstrate that, when lack of smoothness occurs, TTGDA and TTSGDA may slow down, but the nonasymptotic complexity bounds will still be attained as long as certain Lipschitz continuity property is assumed and adaptive stepsizes are carefully chosen (i.e., our algorithms use $(\eta_{\mathbf{x}}^{t},\eta_{\mathbf{y}}^{t})_{t\geq 0}$ instead of $(\eta_{\mathbf{x}},\eta_{\mathbf{y}})$).
@@
Our proof technique in this paper is worth being highlighted. It is different from existing ones for analyzing nested-loop algorithms based on Nouiehed et al. (2019) and Jin et al. (2020). Indeed, the previous frameworks are based on inexact gradient descent algorithms for minimizing a function $\Phi(\cdot)=\max_{\mathbf{y}\in\mathcal{Y}}f(\cdot,\mathbf{y})$, where a subroutine is designed for the inner problem of $\max_{\mathbf{y}\in\mathcal{Y}}f(\mathbf{x},\mathbf{y})$ and which requires $\mathbf{y}_{t}$ to be close to $\mathbf{y}^{\star}(\mathbf{x}_{t})=\mathop{\rm argmax}_{\mathbf{y}\in\mathcal{Y}}f(\mathbf{x}_{t},\mathbf{y})$ at each iteration. Applying this approach to analyze TTGDA and TTSGDA is challenging since $\mathbf{y}_{t}$ might not be close enough to $\mathbf{y}^{\star}(\mathbf{x}_{t})$, so that $\nabla_{\mathbf{x}}f(\mathbf{x}_{t},\mathbf{y}_{t})$ does not approximate a descent direction well. In order to overcome the above issues, we have developed a new proof technique that directly analyzes the concave optimization problem with a slowly changing objective function.
@@
Subsequent to our ICML work, several new algorithms were developed for solving smooth and nonconvex-(strongly)-concave minimax optimization problems (Lin et al., 2020b; Yang et al., 2020b; Ostrovskii et al., 2021b; Yan et al., 2020; Zhang et al., 2020b, 2021; Li et al., 2021; Yang et al., 2022b; Zhao, 2023; Boţ and Böhm, 2023; Xu et al., 2023). In the smooth and nonconvex-strongly-concave setting, a lower bound has been derived for gradient-based algorithms (Zhang et al., 2021; Li et al., 2021) and there exist various nested-loop near-optimal deterministic algorithms (Lin et al., 2020b; Ostrovskii et al., 2021b; Kong and Monteiro, 2021; Zhang et al., 2021). In the smooth and nonconvex-concave setting, no lower bound is known, and the best-known upper bound has been achieved by various nested-loop algorithms (Thekumparampil et al., 2019; Lin et al., 2020b; Yang et al., 2020b; Ostrovskii et al., 2021b; Kong and Monteiro, 2021; Zhao, 2023). Compared to deterministic counterparts, investigations of stochastic algorithms are scarce; indeed, there exists a gap between the best-known lower bound (Li et al., 2021) and upper bound (Yan et al., 2020).
@@
A growing body of research has focused on single-loop algorithms for nonconvex minimax optimization problems (Zhang et al., 2020b; Yang et al., 2022b; Boţ and Böhm, 2023; Xu et al., 2023). For example, Xu et al. (2023) proposed an alternating GDA which converges in both nonconvex-concave and convex-nonconcave settings and showed that the required number of (stochastic) gradient evaluations to find an $\epsilon$-stationary point of $f$ is the best among all single-loop algorithms. Boţ and Böhm (2023) studied the alternating TTGDA and TTSGDA in a more general setting where $f(\cdot,\mathbf{y})$ is Lipschitz and weakly convex but $f(\mathbf{x},\cdot)$ is smooth. They showed that the algorithms can achieve the same complexity bound in this setting as in the smooth setting. Both Zhang et al. (2020b) and Yang et al. (2022b) proposed single-loop algorithms for solving other structured nonconvex-concave minimax optimization problems with theoretical guarantee.
@@ 2 Related Work
Historically, an early concrete instantiation of minimax optimization involved computing a pair of probability vectors $\left(\mathbf{x},\mathbf{y}\right)$, or solving the problem of $\min_{\mathbf{x}\in\Delta^{m}}\max_{\mathbf{y}\in\Delta^{n}}\mathbf{x}^{\top}A\mathbf{y}$ for $A\in\mathbb{R}^{m\times n}$ and simplices $\Delta^{m}\subseteq\mathbb{R}^{m}$ and $\Delta^{n}\subseteq\mathbb{R}^{n}$. This bilinear minimax problem together with von Neumann’s minimax theorem (Neumann, 1928) was a cornerstone in the development of game theory. A simple and generic algorithm was developed for solving this problem where the min and max players implemented a learning procedure in tandem (Robinson, 1951). Subsequently, Sion (1958) generalized von Neumann’s result from bilinear games to general convex-concave games, i.e., $\min_{\mathbf{x}}\max_{\mathbf{y}}f(\mathbf{x},\mathbf{y})=\max_{\mathbf{y}}\min_{\mathbf{x}}f(\mathbf{x},\mathbf{y})$, and triggered a line of research on algorithmic design for convex-concave minimax optimization in continuous time (Kose, 1956; Cherukuri et al., 2017) and discrete time (Uzawa, 1958; Golshtein, 1974; Korpelevich, 1976; Nemirovski, 2004; Auslender and Teboulle, 2009; Nedić and Ozdaglar, 2009; Liang and Stokes, 2019; Mokhtari et al., 2020a, b; Azizian et al., 2020). Two important results about single-timescale GDA are, with diminishing learning rates, they can find an $\epsilon$-approximate saddle point within $O(\epsilon^{-2})$ iterations in the convex-concave setting regardless of smoothness (Auslender and Teboulle, 2009), and $O(\kappa^{2}\log(1/\epsilon))$ iterations in the strongly-convex-strongly-concave setting (Liang and Stokes, 2019).
@@
Smooth and nonconvex-concave minimax optimization problems are a class of computationally tractable problems in the form of Eq. (1.1) which have emerged as a focus in the optimization and machine learning communities (Sinha et al., 2018; Rafique et al., 2021; Sanjabi et al., 2018; Grnarova et al., 2018; Lu et al., 2020; Nouiehed et al., 2019; Thekumparampil et al., 2019; Ostrovskii et al., 2021b; Kong and Monteiro, 2021; Zhao, 2023). Among the existing works, we highlight the work of Grnarova et al. (2018), who proposed a variant of GDA for this class of nonconvex minimax optimization problems and the works of Sinha et al. (2018) and Sanjabi et al. (2018), who studied inexact SGD algorithms (i.e., variants of SGDmax) and proved the convergence guarantee in terms of iteration numbers. Subsequently, Jin et al. (2020) proved the refined convergence results for GDmax and SGDmax in terms of (stochastic) gradient evaluations.
@@
There are some other algorithms developed for solving smooth and nonconvex-concave minimax optimization problems. For example, Rafique et al. (2021) proposed two stochastic algorithms and prove that they efficiently find one stationary point of $\Phi(\cdot)=\max_{\mathbf{y}\in\mathcal{Y}}f(\cdot,\mathbf{y})$. However, their algorithms are nested-loop and the convergence guarantees were established only when $f(\mathbf{x},\cdot)$ is linear and $\mathcal{Y}=\mathbb{R}^{n}$ (see Rafique et al. (2021, Assumption 2)). Another work of Nouiehed et al. (2019) has designed a multistep variant of GDA by incorporating the momentum terms into the scheme [^1] Nouiehed et al. (2019) considered a setting where (i) $\mathbf{x}$ is constrained to a convex set $\mathcal{X}$; (ii) $\mathcal{Y}=\mathbb{R}^{n}$; and (iii) $f$ is smooth in $\mathbf{x}$ over $\mathcal{X}$ and satisfies the Polyak-Łojasiewicz condition in $\mathbf{y}$. Their setting is not more general than ours in all dimensions since $\mathcal{Y}=\mathbb{R}^{n}$ is assumed. Nonetheless, we confirm that their analysis and results are valid in our setting.. In this context, the best-known gradient complexity is $\tilde{O}(\epsilon^{-3})$ and has been achieved by the ProxDIAG algorithm (Thekumparampil et al., 2019). This complexity result was also achieved by an accelerated inexact proximal point algorithm. Indeed, Kong and Monteiro (2021) showed that their algorithm finds one $\epsilon$-stationary point of $\Phi$ within $O(\epsilon^{-3})$ iterations [^2] Kong and Monteiro (2021) considered a more general setting than ours, where $\mathbf{x}$ is constrained to a convex set $\mathcal{X}$ and $f$ is smooth in $\mathbf{x}$ over $\mathcal{X}$. Their analysis and results are valid in our setting.. Solving the subproblem at each iteration is equivalent to minimizing a smooth and strongly convex function and can be approximately done using $\tilde{O}(1)$ gradient evaluations. Combining these two facts yields the desired claim. Subsequent to our ICML work, such best-known complexity bound was also achieved by other algorithms (Lin et al., 2020b; Yang et al., 2020b; Ostrovskii et al., 2021b; Zhao, 2023). Despite appealing theoretical convergence guarantee, all above algorithms are nested-loop and thus can be complex to implement. One would like to understand whether the nested-loop structure is necessary or whether a single-loop algorithm provably converges in smooth and nonconvex-concave settings. An example of such attempts is Lu et al. (2020), where a single-loop algorithm was developed for solving smooth and nonconvex-concave minimax optimization problems. However, their analysis requires some restrictive conditions; e.g., that $f(\cdot,\cdot)$ is lower bounded. In contrast, we only require that $\max_{\mathbf{y}\in\mathcal{Y}}f(\cdot,\mathbf{y})$ is lower bounded. An example which meets our conditions rather than theirs is $\min_{\mathbf{x}\in\mathbb{R}}\max_{\mathbf{y}\in[-1,1]}\mathbf{x}^{\top}\mathbf{y}$. Our less-restrictive conditions make the problem more challenging and our analysis accordingly differs significantly from theirs.
@@
To our knowledge, we are the first to analyze TTGDA and TTSGDA in the nonsmooth and nonconvex-concave setting where $f$ is Lipschitz and $f(\cdot,\mathbf{y})$ is weakly convex. As already noted, Rafique et al. (2021) considered the nonsmooth setting but focused on nested-loop algorithms. Barazandeh and Razaviyayn (2020) and Huang et al. (2021) proposed to study the composite nonconvex-concave setting with nonsmooth regularization terms which are however structured such that the associated proximal mappings can be computed efficiently. The closest work to ours is Boţ and Böhm (2023) who studied the alternating TTGDA and TTSGDA in the nonsmooth and nonconvex-concave setting. However, their setting is still different from ours — they assume that $f(\cdot,\mathbf{y})$ is Lipschitz and weakly convex but $f(\mathbf{x},\cdot)$ is smooth. Their results demonstrate that the algorithms can achieve the same complexity bound in this setting as in the smooth setting. In contrast, as we will show in this paper, relaxing the smoothness of $f$ makes the problem substantially more difficult, especially for deterministic algorithms.

### v2

@@ Abstract
We provide a unified analysis of two-timescale gradient descent ascent (TTGDA) for solving structured nonconvex minimax optimization problems in the form of $\min_{\mathbf{x}}\max_{\mathbf{y}\in\mathcal{Y}}f(\mathbf{x},\mathbf{y})$, where the objective function $f(\mathbf{x},\mathbf{y})$ is nonconvex in $\mathbf{x}$ and concave in $\mathbf{y}$, and the constraint set $\mathcal{Y}\subseteq\mathbb{R}^{n}$ is convex and bounded. In the convex-concave setting, the single-timescale gradient descent ascent (GDA) algorithm is widely used in applications and has been shown to have strong convergence guarantees. In more general settings, however, it can fail to converge. Our contribution is to design TTGDA algorithms that are effective beyond the convex-concave setting, efficiently finding a stationary point of the function $\Phi(\cdot):=\max_{\mathbf{y}\in\mathcal{Y}}f(\cdot,\mathbf{y})$. We also establish theoretical bounds on the complexity of solving both smooth and nonsmooth nonconvex-concave minimax optimization problems. To the best of our knowledge, this is the first systematic analysis of TTGDA for nonconvex minimax optimization, shedding light on its superior performance in training generative adversarial networks (GANs) and in other real-world application problems.
@@ 1 Introduction
In the wake of von Neumann’s seminal work (Neumann, 1928), the problem of finding a saddle-point solution to minimax optimization problems, under various assumptions on the objective functions, has been a major focus of research in many areas, including economics, control theory, operations research and computer science (Basar and Olsder, 1999; Nisan et al., 2007; Von Neumann and Morgenstern, 2007). Recently, minimax optimization theory has begun to see applications in machine learning, with examples including generative adversarial networks (GANs) (Goodfellow et al., 2014), robust statistics (Xu et al., 2009; Abadeh et al., 2015), online adversarial learning (Cesa-Bianchi and Lugosi, 2006), robust training of deep neural networks (Sinha et al., 2018) and distributed computing over networks (Shamma, 2008; Mateos et al., 2010). Current trends in multi-agent learning with large-scale economic systems seem likely to further increase the need for efficient algorithms for computing minimax optima and related equilibrium solutions.
@@
Formally, we study the following minimax optimization problem:
@@
where $f:\mathbb{R}^{m}\times\mathbb{R}^{n}\mapsto\mathbb{R}$ is the objective function and $\mathcal{Y}$ is a constraint set. One of simplest methods for solving the minimax optimization problems is to generalize gradient descent (GD) so that each loop comprises both a descent step and an ascent step. The algorithm, referred to as gradient descent ascent (GDA), performs one-step gradient descent over the min variable $\mathbf{x}$ with a learning rate $\eta_{\mathbf{x}}>0$ and one-step gradient ascent over the max variable $\mathbf{y}$ with another learning rate $\eta_{\mathbf{y}}>0$.
On the positive side, when $f$ is convex in $\mathbf{x}$ and concave in $\mathbf{y}$, there is a vast literature establishing asymptotic and nonasymptotic convergence for average iterates generated by classical single-timescale GDA (i.e., $\eta_{\mathbf{x}}=\eta_{\mathbf{y}}$(see, e.g., Korpelevich, 1976; Chen and Rockafellar, 1997; Nemirovski, 2004; Auslender and Teboulle, 2009; Nedić and Ozdaglar, 2009). In addition, local linear convergence can be established under an additional assumption that $f$ is locally strongly convex in $\mathbf{x}$ and locally strongly concave in $\mathbf{y}$ (Cherukuri et al., 2017; Adolphs et al., 2019). In contrast, there has been no shortage of negative results that highlight the fact that in the general settings, the single-timescale GDA algorithm might converge to limit cycles or even diverge (Mertikopoulos et al., 2018).
@@
We provide detailed pseudocode for TTGDA and its stochastic counterpart (TTSGDA) in Algorithms 1 and 2. Note that Algorithm 2 repeatedly samples while remaining at the current iterate and the subroutine $(\hat{\bf g}_{\mathbf{x}}^{t},\hat{\bf g}_{\mathbf{y}}^{t})=\textsf{SG}(G,\{\xi_{i}^{t}\}_{i=1}^{M},\mathbf{x}_{t},\mathbf{y}_{t})$ can be implemented in general by the updates
$\hat{\bf g}_{\mathbf{x}}^{t}=\tfrac{1}{M}\left(\sum_{i=1}^{M}G_{\mathbf{x}}(\mathbf{x}_{t},\mathbf{y}_{t},\xi_{i}^{t})\right),\quad\hat{\bf g}_{\mathbf{y}}^{t}=\tfrac{1}{M}\left(\sum_{i=1}^{M}G_{\mathbf{y}}(\mathbf{x}_{t},\mathbf{y}_{t},\xi_{i}^{t})\right),$ (1.2)
@@
Our main results in this paper are presented in four theorems. In particular, the first and second theorems (Theorems 4.4 and 4.6) provide a complete characterization of TTGDA and TTSGDA when applied to solve smooth minimax optimization problems. In the nonconvex-strongly-concave setting, TTGDA and TTSGDA require $O(\kappa^{2}\epsilon^{-2})$ and $O(\kappa^{3}\epsilon^{-4})$ (stochastic) gradient evaluations to find an $\epsilon$-stationary point (cf. Definition 3.3) where $\kappa$ is a condition number. In the nonconvex-concave setting, TTGDA and TTSGDA require $O(\epsilon^{-6})$ and $O(\epsilon^{-8})$ (stochastic) gradient evaluations to find an $\epsilon$-stationary point (cf. Definition 3.8). While these results have appeared in our preliminary work (Lin et al., 2020a), the analysis that we provide here is simpler and more direct.
@@
Another two theorems (Theorems 5.4 and 5.6) are the more substantive contributions of this paper. These theorems show that TTGDA and TTSGDA can efficiently find an $\epsilon$-stationary point (cf. Definition 3.8) even when applied to solve nonsmooth minimax optimization problems. The complexity bounds of $\tilde{O}(\epsilon^{-6})$ and $O(\epsilon^{-8})$ in terms of (stochastic) gradient evaluations can be established for the nonconvex-strongly-concave and nonconvex-concave settings. These results are useful since the aforementioned machine learning applications often have the nonsmooth objective functions (e.g., the use of ReLU neural network in GANs). Our theorems demonstrate that, when lack of smoothness occurs, TTGDA and TTSGDA may slow down, but the nonasymptotic complexity bounds will still be attained as long as a certain Lipschitz continuity property is retained and adaptive stepsizes are chosen carefully (i.e., our algorithms use $(\eta_{\mathbf{x}}^{t},\eta_{\mathbf{y}}^{t})_{t\geq 0}$ instead of $(\eta_{\mathbf{x}},\eta_{\mathbf{y}})$).
@@
Our proof technique in this paper is worth being highlighted. It is different from existing techniques for analyzing nested-loop algorithms as exemplified in Nouiehed et al. (2019) and Jin et al. (2020). Indeed, the previous frameworks are based on inexact gradient descent algorithms for minimizing a function $\Phi(\cdot)=\max_{\mathbf{y}\in\mathcal{Y}}f(\cdot,\mathbf{y})$, where a subroutine is designed for the inner problem of $\max_{\mathbf{y}\in\mathcal{Y}}f(\mathbf{x},\mathbf{y})$ and which requires $\mathbf{y}_{t}$ to be close to $\mathbf{y}^{\star}(\mathbf{x}_{t})=\mathop{\rm argmax}_{\mathbf{y}\in\mathcal{Y}}f(\mathbf{x}_{t},\mathbf{y})$ at each iteration. Applying this approach to analyze TTGDA and TTSGDA is challenging since $\mathbf{y}_{t}$ might not be close enough to $\mathbf{y}^{\star}(\mathbf{x}_{t})$, so that $\nabla_{\mathbf{x}}f(\mathbf{x}_{t},\mathbf{y}_{t})$ does not approximate a descent direction well. In order to overcome this issue, we have developed a new proof technique that directly analyzes the concave optimization problem with a slowly changing objective function.
@@
Subsequent to our work in Lin et al. (2020a), several new algorithms were developed for solving smooth and nonconvex-(strongly)-concave minimax optimization problems (Lin et al., 2020b; Yang et al., 2020b; Ostrovskii et al., 2021b; Yan et al., 2020; Zhang et al., 2020b, 2021; Li et al., 2021; Yang et al., 2022b; Zhao, 2023; Boţ and Böhm, 2023; Xu et al., 2023). In the smooth and nonconvex-strongly-concave setting, a lower bound has been derived for gradient-based algorithms (Zhang et al., 2021; Li et al., 2021) and there exist various nested-loop near-optimal deterministic algorithms (Lin et al., 2020b; Ostrovskii et al., 2021b; Kong and Monteiro, 2021; Zhang et al., 2021). In the smooth and nonconvex-concave setting, no lower bound is known, and the best-known upper bound has been achieved by various nested-loop algorithms (Thekumparampil et al., 2019; Lin et al., 2020b; Yang et al., 2020b; Ostrovskii et al., 2021b; Kong and Monteiro, 2021; Zhao, 2023). Compared to deterministic counterparts, investigations of stochastic algorithms are scarce; indeed, there exists a gap between the best-known lower bound (Li et al., 2021) and upper bound (Yan et al., 2020).
@@
A growing body of research has focused on single-loop algorithms for nonconvex minimax optimization problems (Zhang et al., 2020b; Yang et al., 2022b; Boţ and Böhm, 2023; Xu et al., 2023). For example, Xu et al. (2023) proposed an alternating GDA which converges in both nonconvex-concave and convex-nonconcave settings and showed that the required number of (stochastic) gradient evaluations to find an $\epsilon$-stationary point of $f$ is the best among all single-loop algorithms. Boţ and Böhm (2023) studied the alternating TTGDA and TTSGDA in a more general setting where $f(\cdot,\mathbf{y})$ is Lipschitz and weakly convex but $f(\mathbf{x},\cdot)$ is smooth. They showed that the algorithms can achieve the same complexity bound in this setting as in the smooth setting. These two papers extended our proof techniques to analyze the alternating TTGDA and TTSGDA but still assume the smoothness of $f(\mathbf{x},\cdot)$. By contrast, the results in the current paper are the first to demonstrate the efficiency of TTGDA and TTSGDA in the nonsmooth setting. Both Zhang et al. (2020b) and Yang et al. (2022b) proposed single-loop algorithms for solving other structured nonconvex-concave minimax optimization problems with theoretical guarantee.
@@ 2 Related Work
Historically, an early concrete instantiation of minimax optimization involved computing a pair of probability vectors $\left(\mathbf{x},\mathbf{y}\right)$ for solving the problem $\min_{\mathbf{x}\in\Delta^{m}}\max_{\mathbf{y}\in\Delta^{n}}\mathbf{x}^{\top}A\mathbf{y}$ for $A\in\mathbb{R}^{m\times n}$ and simplices $\Delta^{m}\subseteq\mathbb{R}^{m}$ and $\Delta^{n}\subseteq\mathbb{R}^{n}$. This bilinear minimax problem together with von Neumann’s minimax theorem (Neumann, 1928) were a cornerstone in the development of game theory. A simple and generic algorithm was developed for solving this problem where the min and max players implemented a learning procedure in tandem (Robinson, 1951). Subsequently, Sion (1958) generalized von Neumann’s result from bilinear games to general convex-concave games, i.e., $\min_{\mathbf{x}}\max_{\mathbf{y}}f(\mathbf{x},\mathbf{y})=\max_{\mathbf{y}}\min_{\mathbf{x}}f(\mathbf{x},\mathbf{y})$, and triggered a line of research on algorithmic design for convex-concave minimax optimization in continuous time (Kose, 1956; Cherukuri et al., 2017) and discrete time (Uzawa, 1958; Golshtein, 1974; Korpelevich, 1976; Nemirovski, 2004; Auslender and Teboulle, 2009; Nedić and Ozdaglar, 2009; Liang and Stokes, 2019; Mokhtari et al., 2020a, b; Azizian et al., 2020). This line of research has yielded two important conclusions: single-timescale GDA, with decreasing learning rates, can find an $\epsilon$-approximate saddle point within $O(\epsilon^{-2})$ iterations in the convex-concave setting (Auslender and Teboulle, 2009), and $O(\kappa^{2}\log(1/\epsilon))$ iterations in the strongly-convex-strongly-concave setting (Liang and Stokes, 2019).
@@
Smooth and nonconvex-concave minimax optimization problems are a class of computationally tractable problems in the form of Eq. (1.1). They have emerged as a focus in the optimization and machine learning communities (Sinha et al., 2018; Rafique et al., 2021; Sanjabi et al., 2018; Grnarova et al., 2018; Lu et al., 2020; Nouiehed et al., 2019; Thekumparampil et al., 2019; Ostrovskii et al., 2021b; Kong and Monteiro, 2021; Zhao, 2023). Among the existing works, we highlight the work of Grnarova et al. (2018), who proposed a variant of GDA for this class of nonconvex minimax optimization problems and the works of Sinha et al. (2018) and Sanjabi et al. (2018), who studied inexact SGD algorithms (i.e., variants of SGDmax) and established a convergence guarantee in terms of iteration numbers. Subsequently, Jin et al. (2020) proved the refined convergence results for GDmax and SGDmax in terms of (stochastic) gradient evaluations.
@@
There are some other algorithms developed for solving smooth and nonconvex-concave minimax optimization problems. For example, Rafique et al. (2021) proposed two stochastic algorithms and prove that they efficiently find a stationary point of $\Phi(\cdot)=\max_{\mathbf{y}\in\mathcal{Y}}f(\cdot,\mathbf{y})$. However, these algorithms are nested-loop algorithms and the convergence guarantees hold only when $f(\mathbf{x},\cdot)$ is linear and $\mathcal{Y}=\mathbb{R}^{n}$ (see Rafique et al. (2021, Assumption 2)). Nouiehed et al. (2019) have designed a multistep variant of GDA by incorporating momentum terms into the scheme. [^1] Nouiehed et al. (2019) considered a setting where (i) $\mathbf{x}$ is constrained to a convex set $\mathcal{X}$; (ii) $\mathcal{Y}=\mathbb{R}^{n}$; and (iii) $f$ is smooth in $\mathbf{x}$ over $\mathcal{X}$ and satisfies the Polyak-Łojasiewicz condition in $\mathbf{y}$. In this context, the best-known gradient complexity is $\tilde{O}(\epsilon^{-3})$ and has been achieved by the ProxDIAG algorithm (Thekumparampil et al., 2019). This complexity result was also achieved by an accelerated inexact proximal point algorithm. Indeed, Kong and Monteiro (2021) showed that their algorithm finds one $\epsilon$-stationary point of $\Phi$ within $O(\epsilon^{-3})$ iterations. [^2] Kong and Monteiro (2021) considered a more general setting than ours, where $\mathbf{x}$ is constrained to a convex set $\mathcal{X}$ and $f$ is smooth in $\mathbf{x}$ over $\mathcal{X}$. Their analysis and results are valid in our setting. Solving the subproblem at each iteration is equivalent to minimizing a smooth and strongly convex function and can be approximately done using $\tilde{O}(1)$ gradient evaluations. Combining these two facts yields the desired claim. Subsequent to Lin et al. (2020a), this complexity bound was also achieved by other algorithms (Lin et al., 2020b; Yang et al., 2020b; Ostrovskii et al., 2021b; Zhao, 2023). Despite the appealing theoretical convergence guarantee, the above algorithms are all nested-loop algorithms and thus can be complex to implement. One would like to understand whether the nested-loop structure is necessary or whether a single-loop algorithm provably converges in smooth and nonconvex-concave settings. An example of such an investigation is Lu et al. (2020), where a single-loop algorithm was developed for solving smooth and nonconvex-concave minimax optimization problems. However, their analysis requires some restrictive conditions; e.g., that $f(\cdot,\cdot)$ is lower bounded. In contrast, we only require that $\max_{\mathbf{y}\in\mathcal{Y}}f(\cdot,\mathbf{y})$ is lower bounded. An example which meets our conditions rather than theirs is $\min_{\mathbf{x}\in\mathbb{R}}\max_{\mathbf{y}\in[-1,1]}\mathbf{x}^{\top}\mathbf{y}$. Our less-restrictive conditions make the problem more challenging and our analysis accordingly differs significantly from theirs.
@@
To our knowledge, we are the first to analyze TTGDA and TTSGDA in the nonsmooth and nonconvex-concave setting where $f$ is Lipschitz and $f(\cdot,\mathbf{y})$ is weakly convex. As already noted, Rafique et al. (2021) considered the nonsmooth setting but focused on nested-loop algorithms. Barazandeh and Razaviyayn (2020) and Huang et al. (2021) proposed to study the composite nonconvex-concave setting with nonsmooth regularization terms which are however structured such that the associated proximal mappings can be computed efficiently. The closest work to ours is Boţ and Böhm (2023) who studied the alternating TTGDA and TTSGDA in the nonsmooth and nonconvex-concave setting. However, their setting is different from ours—they assume that $f(\cdot,\mathbf{y})$ is Lipschitz and weakly convex but $f(\mathbf{x},\cdot)$ is smooth. Their results demonstrate that the algorithms can achieve the same complexity bound in this setting as in the smooth setting. In contrast, as we will show in this paper, relaxing the smoothness of $f$ makes the problem substantially more difficult, especially for deterministic algorithms.